Waterways is thinking of mass-producing one of its special-order sprinklers. To do so would increase variable costs for all sprinklers by an average of $0.70 per unit. The company also estimates that this change could increase the overall number of sprinklers sold by 10%, and the average sales price would increase $0.20 per unit, waterways currently sells 496,000 sprinkler units at an average selling price of $28.80. The manufacturing costs are $8,039,000 variable and $2,100,879 fixed. Selling and administrative costs are $2,674,600 variable and $793,460 fixed. conne to 5% and le Places. е Round ratio to O decina If Waterways begins mass-producing its special order sprinklers, ho decimal places, e.g. 2,520.) would this affect the company? Effect New Current by Contribution margin ratio by $ Net income LINK TO TEXT LINK TO TEXT LINK TO TEXT Waterways is thinking of mass-producing one of its special-order sprinklers. To do so would increase variable costs for all sprinklers by an average of $0.70 per unit. The company also estimates that this change could increase the overall number of sprinklers sold by 10%, and the average sales price would increase $0.20 per unit. Waterways currently sells 496,000 sprinkler units at an average selling price of $28.80. The manufacturing costs are $8,039,000 variable and $2,100,879 fixed. Selling and administrative costs are $2,674,600 variable and $793,460 fixed. If the average sales price per sprinkler unit did not increase when the company began mass-producing the special-order sprinkler, what would be the effect on the company? (Round answers to 0 decimal places, eg. 590 or 2,520.) by Contribution margin ratio by 5 Profit Open Show Work Click if you would like to Show Work for this question:
Answer #1

5A. Current

Variable cost per unit = 10713600/496000

= $21.60.

Particulars Amount $
Total Sales (496000*28.80) 14284800
Less: Total Variable Costs (8039000+2674600) 10713600
Contribution Margin 3571200
Less: Total Fixed Costs (2100879+793460) 2894339
Net Income 676861

Contribution Margin Ratio = Contribution Margin/Sales * 100

= 3571200/14284800 * 100

= 25%

New

Selling Price = 28.80+0.20

= $29.00

Variable Cost = 21.60 + 0.70

= $22.30

Number of units estimated to be sold = 496000 * 110%

= 545600

Particulars Amount $
Total Sales (545600*29) 15822400
Less: Total Variable Costs (545600*22.30) 12166880
Contribution Margin 3655520
Less: Total Fixed Costs (2100879+793460) 2894339
Net Income 761181

Contribution Margin Ratio = Contribution Margin/Sales * 100

= 3655520/15822400 * 100

= 23.10%

Thus,

Particulars Current New Effect Effect
Contribution Margin Ratio 25% 23.10% Decrease 1.90%(directly) or 7.6%(relatively)
Net Income $676861 $761181 Increase $84,320

5B.

New

Selling Price = 28.80

Variable Cost = 21.60 + 0.70

= $22.30

Number of units estimated to be sold = 496000 * 110%

= 545600

Particulars Amount $
Total Sales (545600*28.80) 15713280
Less: Total Variable Costs (545600*22.30) 12166880
Contribution Margin 3546400
Less: Total Fixed Costs (2100879+793460) 2894339
Net Income 652061

Contribution Margin Ratio = Contribution Margin/Sales * 100

= 3546400/15713280 * 100

= 22.57%

Thus,

Particulars Current New Effect Effect
Contribution Margin Ratio 25% 22.57% Decrease 2.43%(directly) or 9.72%(relatively)
Net Income $676861 $652061 Decrease $24,800